OpenStudy (anonymous):

A sample of gas occupies a volume of 482 mL at a temperature of 282 K. At what temperature will this sample of gas occupy a volume of 502 mL? Show all of the work used to solve this problem.

OpenStudy (anonymous):

Use V1/T1 = V2/T2

OpenStudy (anonymous):

its not set up for it yet. Wouldn't i have to use PV=nRT first? to find the moles then use V1/T1=V2/T2?

OpenStudy (anonymous):

dont need to can use this directly its called Boyle's Law

OpenStudy (anonymous):

Would you show me how to? Please?

OpenStudy (lgbasallote):

\[PV = nRT\] since pressure and moles are constant you can just cancel them \[V = RT\] \[\frac{V_1}{T_1} = R\] \[\frac{482}{282} = R\] \[R = 1.709\] so now solve for T_2 \[\frac{V_2}{T_2} = 1.709\] \[\frac{V_2}{1.709} = T_2\] \[\frac{502}{1.709} = T_2\] \[293.70 K = T_2\] that's if you really want to use PV = nRT but boyle's law just states that when pressure and moles are constant then \[\frac{V_1}{T_1} = \frac{V_2}{T_2}\]

OpenStudy (anonymous):

Thank you so much both you. :) I appreciate it.